# Design of a Wearable Parallel Electrical Impedance Imaging System for Healthcare

### v1

@@ 2.3 Multiplexer and Voltage Measurement
Channel switching is implemented using the ADG706 multiplexer. The ADG706 is controlled via four binary address lines (A0, A1, A2, and A3) to select one of the 16 input terminals (S1–S16) and connect it to the common output terminal D. Compared with traditional relay-based switching methods, the use of the ADG706 may introduce parasitic capacitance effects. Nevertheless, it is still selected as the channel switching device due to its smaller size and faster switching speed.
In EIT, a constant-amplitude current signal is commonly used for excitation. In this context, the impedance of the region under test directly influences the boundary voltage. Specifically, a higher impedance results in a larger boundary voltage, while a lower impedance leads to a smaller one. To ensure that the boundary voltage remains within the ADC’s sampling range, a programmable gain amplifier (PGA) is typically employed to adjust the signal and prevent it from exceeding the input limits. However, when the boundary voltage is very low, the PGA must apply a high gain, which may also amplify noise, thereby degrading the system’s signal-to-noise ratio (SNR). In this study, the adjacent excitation-adjacent measurement mode is adopted, using a constant-amplitude AC voltage to excite two adjacent electrodes among the 16 electrodes. In this mode, the maximum boundary voltage difference between the two adjacent electrodes is equal to the amplitude of the excitation voltage. Since the boundary voltage difference has a relatively large amplitude and remains within the ADC’s maximum input range, it can be ensured that the boundary voltage stays within a suitable range for sampling. Therefore, with fixed-amplitude voltage excitation, a PGA is no longer required to adjust the boundary voltage. Instead, voltage acquisition can be accomplished using only a voltage follower circuit.
When using a four-terminal semi-parallel configuration for impedance acquisition and selecting two electrodes for voltage excitation, the four voltage detection circuits are connected to four electrodes via four 16-to-1 multiplexer chips. In this configuration, if the 16 input terminals (S1-S16) of each multiplexer chip are directly connected to the 16 electrodes, each voltage detection electrode introduces a parasitic capacitance of 200 pF. This capacitance results from the activation of the input channels (S1-S16) on the multiplexer chips. When the signal frequency exceeds 50 kHz, parasitic capacitance may induce leakage currents, which can distort the current field distribution within the object under test and compromise measurement accuracy. To mitigate this issue, the 16 input terminals of the four multiplexer chips should not be connected directly to the electrodes but instead routed through voltage followers (i.e. buffers), as shown in Fig.9. Owing to their high input impedance and low output impedance, voltage followers effectively suppress leakage currents, thereby ensuring accurate current field distribution and precise measurement results.
@@ 2.4 I2C Multiplexer and External Clock
This paper implements communication with five AD5933 chips through the I2C multiplexer chip TCA9548A. Developed by Texas Instruments (TI), the TCA9548A is capable of splitting a single I2C bus signal into eight independent channels, each of which can be individually enabled or disabled. This approach allows multiple I2C devices with identical addresses to share the same I2C bus, effectively avoiding address conflicts and enabling simultaneous activation of several I2C channels. By appropriately configuring the registers of the TCA9548A, simultaneous transmission of measurement commands to five AD5933 chips was successfully realized, ensuring synchronized data acquisition.
This paper selects the ICS553 as the clock distribution chip. Developed by Renesas, the ICS553 is a clock buffer capable of generating up to four output clock signals from a single reference clock input. Employing an external clock primarily ensures synchronized data acquisition. The ADC sampling frequency of the AD5933 is determined by dividing the external clock frequency by 16. Thus, using a unified external clock allows the ADC sampling frequencies of the five parallel AD5933 chips to remain consistent. Fig.10(a) illustrates the connection between the clock signal outputs of the ICS553 and the clock signal inputs of the AD5933. The purpose of this circuit is to limit the signal amplitude. The equivalent circuit is presented in Fig.10(b).
@@ 2.5.1 Serial Measurement
After the system is powered on, the STM32 writes specific values to the control register of the TCA9548A to connect to the first AD5933 and initialize it.
@@
In the two-terminal serial impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to apply a voltage excitation, while the first AD5933 measures the current. In the four-terminal serial impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to perform a total of 17 voltage excitations. During the first excitation, the first AD5933 measures the current; from the second excitation onward, the first AD5933 sequentially measures the voltage of each electrode until all 16 electrodes have been measured.
@@
During current or voltage measurements, STM32 sends commands to the control register of the AD5933 to put it into standby mode, initialize the start frequency, and begin frequency scanning. STM32 then continuously queries the status register of the AD5933 to confirm whether the Discrete Fourier Transform (DFT) has been completed. If completed, STM32 reads the real and imaginary part data from the AD5933 registers.
@@
The voltage excitation electrodes are then changed via the multiplexer. After each change in the two electrodes used for voltage excitation, STM32 performs an impedance calculation. Once all impedance data collection is completed, STM32 transmits a complete frame of impedance data to the PC for further processing. Fig.11(a) shows the electrode excitation and measurement sequence in the two-terminal serial measurement mode. Fig.11(b) shows the excitation and measurement sequence in the four-terminal serial measurement mode. Fig.13(a) illustrates the program flowchart for both the two-terminal and four-terminal serial measurement modes.
@@
2.5.2 Parallel Measurement
After the system is powered on, STM32 first writes specific values to the control register of the TCA9548A to connect the first, second, third, fourth, and fifth AD5933 devices and complete their initialization.
In the four-terminal half-parallel impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to apply a total of 5 voltage excitations. During each excitation, the first AD5933 measures the current, while the second, third, fourth, and fifth AD5933 devices measure the voltage of one electrode from the 16 electrodes. During the first voltage excitation, the second, third, fourth, and fifth AD5933 devices measure the voltage at electrodes 1, 5, 9, and 13, respectively; during the second excitation, they measure the voltage at electrodes 2, 6, 10, and 14, respectively; during the third excitation, they measure the voltage at electrodes 3, 7, 11, and 15, respectively; during the fourth excitation, they measure the voltage at electrodes 4, 8, 12, and 16, respectively; and during the fifth excitation, they measure the voltage at electrodes 5, 9, 13, and 1, respectively. This ensures that the voltage differences between adjacent electrodes (e.g., 1-2, 2-3, 3-4, 4-5) are measured by the second AD5933, the voltage differences between other adjacent electrodes (e.g., 5-6, 6-7, 7-8, 8-9) are measured by the third AD5933, and so on. This design ensures that the voltage difference between each pair of adjacent electrodes is measured by the same AD5933, which improves the common-mode rejection ratio of the system’s voltage measurements.
During the current or voltage measurement process, STM32 first writes specific values to the TCA9548A register to connect it with the first AD5933, which is responsible for generating the voltage excitation signal. Then, STM32 sends a command to the control register of the first AD5933 to put it into standby mode, initialize the start frequency, and begin frequency scanning. This allows the AD5933 to generate the excitation voltage and measure the excitation current after the voltage passes through the test specimen. Next, STM32 writes new values to the TCA9548A register to connect it with the second, third, fourth, and fifth AD5933 devices, and follows the same steps by sending standby commands, initializing start frequency commands, and starting frequency scanning commands to these AD5933 for voltage measurement. During the Discrete Fourier Transform (DFT), the excitation signal is applied first, followed by the measurement of boundary voltages. At the same time, the frequency of the excitation signal is set to an integer multiple of the DFT data acquisition frequency for 1024 points. This ensures that each 1024-point acquisition window contains a complete number of excitation signal cycles, avoiding measurement errors caused by incomplete waveforms. A specific diagram is shown in Fig.12. The formula for DFT data sampling frequency is:
$f_{DFT}=\frac{f_{ADC}}{n}$ (9)
Where $f_{ADC}$ is the ADC sampling frequency and n is the number of collected samples (such as 1024 samples). It can be calculated that when the clock frequency is 16 MHz, the ADC sampling frequency is 1 MHz, and the DFT data sampling frequency is $\frac{1\,\text{MHz}}{1024}$.
Once the frequency scan begins, STM32 writes specific values to the TCA9548A register to sequentially connect it with the first, second, third, fourth, and fifth AD5933 devices. With each connection, STM32 continuously queries the status register of the AD5933 to confirm whether the Discrete Fourier Transform (DFT) has been completed. If completed, STM32 reads the real and imaginary part data from the AD5933 registers.
The voltage excitation electrodes are then switched by multiplexing. After each switch of the two excitation electrodes, the system calculates the impedance based on one current measurement and 16 voltage measurements from the electrodes, resulting in 16 impedance values. After collecting 256 impedance data points from 16 electrode pairs in the adjacent excitation-adjacent measurement mode, STM32 transmits a complete frame of impedance data to the PC for further processing. Fig.11(c) shows the excitation and measurement sequence in the four-terminal semi-parallel measurement mode. Fig.13(b) illustrates the program flowchart of the four-terminal semi-parallel impedance measurement system.
@@ 2.6 Impedance Calculation
When using the two-terminal measurement method, it is important to understand that the measured data does not directly represent the real and imaginary components of the impedance. Instead, these data reflect the frequency-domain representation of the current signal. Since the excitation voltage is known, the real and imaginary components of the impedance can be calculated using the measured current data in conjunction with a predetermined gain factor.
The phase angle is given by the following formula:
$\phi=\tan^{-1}\left(\frac{Re}{Im}\right)$ (10)
Here, $Re$ represents the real part, and $Im$ represents the imaginary part. The phase angle of the current signal measured by the AD5933 results from various phase shifts introduced by different elements in the measurement system. These include the phase of the excitation voltage generated by the AD5933, the phase shift of the current signal relative to the voltage signal within the sample under test, the phase shift introduced by the I-V conversion circuitry, and the phase shift from the AD5933’s internal low-pass filter. These combined phase shifts contribute to the final measured phase angle.
To measure the phase shift of the current signal relative to the voltage signal in the sample under test, a two-step approach is required. Step 1: Connect a pure resistor to the circuit and measure the phase angle of the resulting signal with the AD5933. The measured phase at this point represents the system phase, which does not include the phase shift of the current signal relative to the voltage signal within the sample. Instead, it includes the excitation voltage phase generated by the AD5933, the phase shift introduced by the I-V conversion circuitry, and the phase shift from the AD5933’s internal low-pass filter. Step 2: Replace the pure resistor with the actual sample under test and measure again. Now, the phase shift of the current signal relative to the voltage signal in the sample can be determined using the following formula: $\phi_{cal}=\phi_{unk}-\phi_{res}$. Here, $\phi_{res}$ is the system phase measured in Step 1 with the known pure resistor, and $\phi_{unk}$ is the total phase measured in Step 2 with the actual sample. By calculating the difference between these two phase measurements, the phase shifts introduced by the system itself can be eliminated, allowing for the calculation of the phase shift of the current signal relative to the voltage signal within the sample.
If: $|I|\cdot e^{j\phi_{unk}}=Re_{unk}+Im_{unk}j$, $|I|\cdot e^{j\phi_{res}}=Re_{res}+Im_{res}j$, where $Re_{unk}$ and $Im_{unk}$ represent the real and imaginary parts of the current signal measured by the AD5933 when the test sample is connected, and $Re_{res}$ and $Im_{res}$ represent the real and imaginary parts of the current signal measured when a pure resistor is connected. According to the phase correction formula $\phi_{cal}=\phi_{unk}-\phi_{res}$, we obtain:
@@
Next, define the complex expression for the calibrated current signal as: $|I|\cdot e^{j\phi_{cal}}=Re_{cal}+Im_{cal}j$, we then have:
$Re_{cal}+Im_{cal}j=\left(Re_{unk}+Im_{unk}j\right)\frac{\left(Re_{res}-Im_{res}j\right)}{\sqrt{Re_{res}^{2}+Im_{res}^{2}}}$ (12)
By solving Eq.12, the following results:
$\begin{cases}Re_{cal}=\frac{Re_{unk}Re_{res}+Im_{unk}Im_{res}}{\sqrt{Re_{res}^{2}+Im_{res}^{2}}}\\ Im_{cal}=\frac{Re_{res}Im_{unk}-Re_{unk}Im_{res}}{\sqrt{Re_{res}^{2}+Im_{res}^{2}}}\end{cases}$ (13)
For a two-terminal system, the complex impedance Z can be calculated as:
@@
Here, $Gain$ is the gain factor, a constant determined during the calibration process. The magnitude of the impedance $|Z|$ is calculated by:
$|Z|=\frac{1}{Gain\cdot\sqrt{Re_{cal}^{2}+Im_{cal}^{2}}}$ (15)
The imaginary part of the impedance, $Z_{Im}$, is calculated as:
$Z_{Im}=\frac{-Im_{cal}}{\text{Gain}\cdot\left(Re_{cal}^{2}+Im_{cal}^{2}\right)}$ (16)
When using the four-terminal measurement method, the voltage signal from one of the 16 electrodes is processed using the Discrete Fourier Transform (DFT), resulting in the following expression: $X_{1}(f)=Re_{elec1}+Im_{elec1}j$, This result is calculated by:
$X_{1}(f)=\textstyle\sum_{n=0}^{N-1}x_{1}(n)\cdot e^{-j\frac{2\pi}{N}kn},\quad k=0,1,2,\dots,N-1$ (17)
where $X_{1}(f)$ is the frequency-domain component at frequency k, $x_{1}(n)$ is the discrete-time signal acquired by the ADC, and $e^{-j\frac{2\pi}{N}kn}$ is the rotation factor. $N$ is the total number of points used in the DFT calculation. DFT satisfies linearity, which means for any constants $a$ and $b$, and any signals $x(n)$ and $y(n)$, the following holds:
${DFT}\{ax(n)+by(n)\}=aX(f)+bY(f)$ (18)
Thus, for the voltage difference signal between two electrodes, $x_{1}(n)-x_{2}(n)$, its DFT can be expressed as:
${DFT}\{x_{1}(n)-x_{2}(n)\}=X_{1}(f)-X_{2}(f)$ (19)
By substituting the DFT results of the voltage signals, $X_{1}(f)=Re_{elec1}+Im_{elec1}j$ and $X_{2}(f)=Re_{elec2}+Im_{elec2}j$, into the above equation, we obtain:
${DFT}\{x_{1}(n)-x_{2}(n)\}$ $=\left(Re_{\text{elec1}}-Re_{\text{elec2}}\right)$
@@
Therefore, the DFT of the voltage difference between two electrodes can be obtained by individually performing DFT on each electrode’s voltage signal and then calculating the differences of their real and imaginary parts.
If the DFT result of the voltage difference signal between the two electrodes is: $X_{diff}(f)=\left(Re_{elec1}-Re_{elec2}\right)+\left(Im_{elec1}-Im_{elec2}\right)j$, and the DFT result of the current signal is: $X_{I}(f)=Re_{I}+Im_{I}j$, the complex impedance $Z(f)$ can be calculated as:
$Z(f)={Gain}\cdot\frac{X_{\text{diff}}(f)}{X_{I}(f)}$ (21)
Similarly, $Gain$ is the gain factor determined during calibration. The magnitude of the complex impedance $|Z(f)|$ is calculated as:
$|Z(f)|=\sqrt{Z_{Re}(f)^{2}+Z_{Im}(f)^{2}}$ (22)
@@ 2.7 Graphical User Interface
PyQt5 allows developers to build graphical user interfaces (GUIs) using Python. In this article, we will use PyQt5 to design and implement the GUI. The system software is composed of six key components: data reading thread, data saving thread, calibration data thread, imaging thread, interface mapping thread, and the main program. With multi-threading support, tasks like real-time data processing and image updates are handled by separate threads, ensuring a smooth and responsive main interface without lag or delays.
Fig.14 illustrates the GUI. This interface not only supports real-time imaging and data visualization but also allows for data saving and calibration. While users analyze the EIT images, they can conveniently save the data for later offline analysis and experiment reproduction. On the left side, a dynamic line chart displays the impedance data collected by the acquisition device. The horizontal axis represents 256 impedance points, while the vertical axis shows their corresponding values. The real-time updating chart provides a clear visualization of the impedance data distribution, helping users monitor changes and detect potential anomalies. On the right side, the EIT image is displayed. The colors in the image correspond to different impedance values, with red indicating low impedance areas and blue indicating high impedance areas. This color differentiation helps users more clearly observe the impedance distribution within the scanned area. At the bottom of the interface, various functional buttons are provided, including ”Start,” ”Pause,” and ”Calibration.” Users can effortlessly start or stop data acquisition, refresh the interface, save data, and control imaging with these buttons. The layout of the buttons is clean and intuitive, greatly enhancing user convenience.

### v2

@@ 2.3 Multiplexer and Voltage Measurement
Channel switching utilizes the ADG706 multiplexer. Although the ADG706 can introduce parasitic capacitance compared to traditional relay-based methods, it remains the preferred option due to its smaller size and faster switching speed.
In EIT, excitation is commonly performed using a constant-amplitude current signal. In this setup, the impedance of the region under test directly influences the resulting boundary voltage. To ensure that this boundary voltage remains within the sampling range of the ADC, a Programmable Gain Amplifier (PGA) is typically employed to dynamically adjust the voltage level. However, in this study, excitation is performed using a constant-amplitude voltage signal. As a result, a PGA is no longer necessary for adjusting the boundary voltage, and voltage acquisition can be achieved using only a voltage follower circuit.
In the four-terminal parallel mode, if the four voltage detection circuits are directly connected to four electrodes via four multiplexer chips, each of these electrodes introduces a parasitic capacitance of 200 pF. This parasitic capacitance results from the activation of the multiplexer channels. At frequencies above 50 kHz, leakage currents induced by parasitic capacitance may alter the current field distribution within the object under test, leading to reduced measurement accuracy. To prevent this, the 16 input terminals of the four multiplexer chips should be isolated from the electrodes using voltage followers (buffers), as depicted in Fig.9.
@@ 2.4 I2C Multiplexer and External Clock
Communication with five AD5933 chips is implemented through the I2C multiplexer chip TCA9548A. Developed by Texas Instruments (TI), the TCA9548A is capable of splitting a single I2C bus signal into eight independent channels, each of which can be individually enabled or disabled. This approach allows multiple I2C devices with identical addresses to share the same I2C bus, effectively avoiding address conflicts and enabling simultaneous activation of several I2C channels. By appropriately configuring the registers of the TCA9548A, simultaneous transmission of measurement commands to five AD5933 chips was successfully realized, ensuring synchronized data acquisition.
The ICS553 is selected as the clock distribution chip. Developed by Renesas, the ICS553 is a clock buffer capable of generating up to four output clock signals from a single reference clock input. Employing an external clock primarily ensures synchronized data acquisition. The ADC sampling frequency of the AD5933 is determined by dividing the external clock frequency by 16. Thus, using a unified external clock allows the ADC sampling frequencies of the five parallel AD5933 chips to remain consistent. Fig.10(a) illustrates the connection between the clock signal outputs of the ICS553 and the clock signal inputs of the AD5933. The purpose of this circuit is to limit the signal amplitude. The equivalent circuit is presented in Fig.10(b).
@@ 2.5.1 Serial Measurement
After the system is powered on, the first AD5933 is connected and initialized. In the two-terminal serial impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to apply a voltage excitation, while the first AD5933 measures the current. In the four-terminal serial impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to perform a total of 17 voltage excitations. During the first excitation, the first AD5933 measures the current; from the second excitation onward, the first AD5933 sequentially measures the voltage of each electrode until all 16 electrodes have been measured.
@@
During current or voltage measurements, first, a command to start frequency scanning is sent to the control register of the AD5933. Then, the status register of the AD5933 is continuously queried to check whether the Discrete Fourier Transform (DFT) has been completed. Once completed, the STM32 reads the real and imaginary part data from the AD5933 registers.
@@
The voltage excitation electrodes are then changed via the multiplexer. After each change in the two electrodes used for voltage excitation, STM32 performs an impedance calculation. Once all impedance data collection is completed, STM32 transmits a complete frame of impedance data to the PC for further processing. Fig.12(a) shows the excitation and measurement sequence in the two-terminal serial measurement mode. Fig.12(b) shows the excitation and measurement sequence in the four-terminal serial measurement mode. Fig.13(a) illustrates the program flowchart for both the two-terminal and four-terminal serial measurement modes.
2.5.2 Parallel Measurement
After the system is powered on, each of the five AD5933 chips is connected and initialized. In the four-terminal parallel impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to perform a total of five voltage excitations. During each excitation, the first AD5933 consistently measures the current, while the second to fifth AD5933 chips sequentially measure the voltages at specific electrodes as follows: in the first excitation, electrodes 1, 5, 9, and 13 are measured; in the second excitation, electrodes 2, 6, 10, and 14; in the third excitation, electrodes 3, 7, 11, and 15; in the fourth excitation, electrodes 4, 8, 12, and 16; and in the fifth excitation, electrodes 5, 9, 13, and 1. This design ensures that the voltage difference between each pair of adjacent electrodes is measured by the same AD5933, thereby improving the common-mode rejection ratio of the system’s voltage measurements.
During the current or voltage measurement process, the system first connects to the first AD5933 and sends a frequency scanning start command, causing it to generate an excitation voltage and measure the current. Subsequently, the system simultaneously connects to the second, third, fourth, and fifth AD5933 chips and sends a frequency scanning start command, enabling these AD5933 chips to perform voltage measurements. The method involves first applying the voltage excitation signal, followed by measuring the boundary voltages. Meanwhile, the frequency of the excitation signal is set to be an integer multiple of the data acquisition frequency for the 1024-point DFT. This ensures that each 1024-point acquisition window captures an integer number of excitation cycles, thereby avoiding measurement errors caused by incomplete waveforms. A detailed illustration is provided in in Fig.11. The formula for the DFT data sampling frequency is:
$f_{DFT}=\frac{f_{ADC}}{n}$ (9)
Where $f_{ADC}$ is the ADC sampling frequency and n is the number of collected samples (such as 1024 samples).
Once the frequency scan begins, STM32 sequentially connects to the first, second, third, fourth, and fifth AD5933 chips. During each connection, it continuously queries the status register of the AD5933 to check whether the DFT has been completed. If completed, STM32 reads the real and imaginary part data from the AD5933 registers.
The voltage excitation electrodes are then changed via the multiplexer. After each change in the two electrodes used for voltage excitation, STM32 performs an impedance calculation. Once all impedance data collection is completed, STM32 transmits a complete frame of impedance data to the PC for further processing. Fig.12(c) shows the excitation and measurement sequence in the four-terminal parallel measurement mode. Fig.13(b) illustrates the program flowchart of the four-terminal parallel impedance measurement mode.
@@ 2.6 Impedance Calculation
In the two-terminal measurement mode, the measured quantity is the current. Given that the excitation voltage is fixed, the real and imaginary parts of the impedance can be derived from the current data together with a known gain factor.
To measure the phase shift of the current signal relative to the voltage signal in the sample under test, a two-step process is required. In the first step, a pure resistor is connected to the circuit for measurement. The phase angle measured by the AD5933 at this point represents the system phase. In the second step, the pure resistor is replaced with the sample under test, and another measurement is taken. The phase angle measured by the AD5933 now represents the total phase. The phase shift of the current signal relative to the voltage signal in the sample can be calculated using the formula: $\phi_{cal}=\phi_{unk}-\phi_{res}$, where $\phi_{res}$ is the system phase and $\phi_{unk}$ is the total phase. By subtracting the two values, the phase shift introduced by the system itself is eliminated, allowing the phase shift of the current signal relative to the voltage signal in the sample to be determined.
If: $|I|\cdot e^{j\phi_{unk}}=Re_{unk}+Im_{unk}j$ and $|I|\cdot e^{j\phi_{res}}=Re_{res}+Im_{res}j$, where $Re_{unk}$ and $Im_{unk}$ represent the real and imaginary parts of the current signal measured by the AD5933 when the test sample is connected, and $Re_{res}$ and $Im_{res}$ represent the real and imaginary parts of the current signal measured when a pure resistor is connected. According to the phase correction formula $\phi_{cal}=\phi_{unk}-\phi_{res}$, the complex representation of the calibrated current signal is:
@@
the complex impedance Z can be calculated as:
@@
$Gain$ is the gain factor, a constant determined during the calibration process.
In the four-terminal measurement mode, the voltage signals $x_{1}(n)$ and $x_{2}(n)$ from two adjacent electrodes are transformed using the DFT. The results are: $X_{1}(f)=Re_{elec1}+Im_{elec1}j$, $X_{2}(f)=Re_{elec2}+Im_{elec2}j$. The DFT satisfies the linearity property, so the DFT result of the voltage difference signal between the two electrodes is:
${DFT}\{x_{1}(n)-x_{2}(n)\}$ $=X_{1}(f)-X_{2}(f)$
$=\left(Re_{\text{elec1}}-Re_{\text{elec2}}\right)$
@@
If the DFT result of the current signal is: $X_{I}(f)=Re_{I}+Im_{I}j$, the complex impedance $Z(f)$ can be calculated as:
$Z(f)={Gain}\cdot\frac{X_{1}(f)-X_{2}(f)}{X_{I}(f)}$ (13)
@@ 2.7 Graphical User Interface
PyQt5 enables developers to build graphical user interfaces (GUIs) using Python. In this work, we use PyQt5 to design and implement the GUI. The system software consists of six main components: the data reading thread, data saving thread, calibration thread, imaging thread, interface mapping thread, and the main program. By introducing multi-threading, operations such as real-time data processing and image updates can be performed in background threads, ensuring the main interface remains smooth and responsive without stuttering or lag.
Fig.14 shows the GUI, which integrates real-time imaging and data visualization, as well as data saving and calibration functions. While intuitively analyzing the EIT images, users can also easily save the data for subsequent offline analysis and experimental reproduction.
On the left side, a dynamic line chart presents the impedance data collected by the acquisition device. The horizontal axis represents 256 impedance points, while the vertical axis shows their corresponding values. This real-time updating chart offers an intuitive view of the impedance distribution, enabling users to track changes and identify potential anomalies. On the right side, the EIT image is displayed. The image uses color to represent different impedance values—red indicates regions of low impedance, while blue indicates high impedance. This visual differentiation helps users clearly observe the impedance distribution within the scanned area. At the bottom of the interface, multiple functional buttons are available, such as ”Start,” ”Pause,” and ”Calibration.” These buttons allow users to easily start or stop data acquisition, refresh the interface, save data, and manage imaging settings. The layout is clean and user-friendly, significantly improving operational efficiency and ease of use.